calculate the average speed for driving 60 mph for 100 miles first, then flying 360 mph for 400 miles.
Answer by robertb(5830)

The first part of the trip was made in 100mi/(60mi/hr) = 5/3 hrs
The 2nd part of the trip was made in 400mi/(360mi/hr) = 10/9 hrs.
Therfore the trip took 5/3hr + 10/9hr = 25/9 hrs.
Total distance traveled was 100 miles + 400miles = 500 miles.
Therefore average speed was 500miles/(25/9 hr) = 180 miles per hour.